12. Consider five spin-1/2 elementary particles (distinguishable and with no external fields present). What is the probability that four have spin up and the other has spin down, and how many different configurations of the five could give this result?
Answer #1

Total different combinations = 2*2*2*2*2 = 25 =32

Combinations with 4 having spin up and other having spin down = 5C4 = 5

Probability that four have spin up while other have spin down

= 5/32 = 0.15625